# Dynamics of Agentic Loops in Large Language Models: A Geometric Theory of Trajectories From Semantic Contraction to Exploratory Divergence

### v3

@@ Abstract
Agentic systems based on large language models operate through recursive feedback loops, where each output becomes the next input in an iterative transformation process. Understanding the geometric dynamics of these agentic loops, whether they converge, diverge or exhibit more complex behaviors, remains poorly characterized. This paper develops a geometric framework for analyzing agentic loop trajectories in semantic embedding spaces, treating iterative transformations as discrete dynamical systems.
We formalize the distinction between the artifact space (where linguistic transformations occur) and the semantic embedding space (where geometric measurements are performed). Since standard cosine similarities exhibit systematic bias due to embedding anisotropy, we introduce calibrated similarity using isotonic regression, which eliminates bias and improves correlation with human semantic judgments while maintaining high local stability. This calibrated metric enables precise geometric analysis through trajectories (temporal sequences of embeddings), clusters (regions of semantic coherence) and attractors (convergence points).
Experimentally, we characterize two fundamental dynamical regimes through singular agentic loops. A contractive loop performing iterative paraphrasing exhibits geometric convergence toward a stable attractor with decreasing dispersion. An exploratory loop alternating summarization and negation shows unbounded divergence with no stable clustering structure. These regimes demonstrate qualitatively distinct trajectory geometries, contraction versus expansion in semantic space.
Prompt design directly controls the dynamical regime, enabling systematic engineering of convergent versus divergent behaviors. This geometric framework provides the first rigorous quantitative approach to analyzing iterative transformation dynamics in LLMs, with applications to trajectory prediction, stability analysis and controlled composition of complex multi-step transformations.
@@ 1 Introduction
Large language models (LLMs) enable autonomous agentic systems that operate through recursive feedback loops, where each model output becomes the next input in an iterative transformation process. These agentic loops appear in diverse applications: iterative refinement systems that progressively improve outputs through self-critique, multi-step reasoning chains that decompose complex problems into intermediate steps and autonomous agents that maintain extended interactions with environments or users. Understanding the dynamics of these loops, whether they converge toward stable states, diverge unboundedly or exhibit complex oscillatory behaviors, remains poorly characterized. Despite their increasing deployment in reasoning, planning and generation tasks, we lack a rigorous framework for analyzing how iterative transformations behave geometrically in semantic space.
Characterizing agentic loop dynamics has direct implications across multiple domains. For system design, knowing whether a loop exhibits convergent or divergent behavior enables engineers to select appropriate prompts and parameters for specific tasks, stable convergence is desirable for refinement and summarization, while controlled divergence may be useful for exploration and generation. For reliability and safety, detecting convergence properties helps identify runaway feedback loops that could produce unbounded drift or oscillation. For theoretical understanding, geometric trajectory analysis provides a quantitative lens for studying how language model transformations behave as discrete dynamical systems. The ability to predict, measure and control these dynamics is fundamental to deploying reliable iterative generative systems.
@@
The main contributions of this work are:
  1. 1.
    A geometric framework for analyzing agentic loops. We distinguish the artifact space (where text transformations occur) from the semantic embedding space (where geometric measurements are performed) and formalize how transformations induce measurable trajectories, temporal sequences of embeddings that trace the loop’s evolution through semantic space.
  2. 2.
    A calibrated similarity metric. To overcome the fundamental obstacle that pretrained embeddings exhibit anisotropy, concentrating cosine similarities around uninformative high values, we develop a calibrated metric using isotonic regression on human semantic judgments. This calibration eliminates bias and improves correlation while maintaining high local stability (98%), enabling precise trajectory analysis.
  3. 3.
    The first systematic empirical characterization of dynamical regimes. Through singular agentic loops, we identify and characterize two fundamental regimes: contractive convergence and exploratory divergence. We demonstrate that prompt design directly controls the geometric behavior of iterative transformations, establishing that agentic loop dynamics are predictable and controllable through prompt engineering.
The remainder of the paper is organized as follows. Section 2 reviews related work on iterative refinement, agentic systems, dynamical systems and embedding geometry. Sections 35 construct the geometric measurement infrastructure: embedding space construction, similarity calibration and stability validation. Section 6 formalizes the theoretical framework, defining agentic loops, trajectories, clusters, attractors and dynamical regimes. Section 7 presents experimental trajectory studies of contractive and exploratory singular loops, revealing fundamentally distinct geometric behaviors. Section 8 discusses implications for composite loops and controlled dynamics. Section 9 concludes.
An earlier version of this work was released as an arXiv preprint.

### v4

# Geometric Dynamics of Agentic Loops in Large Language Models Trajectories, Attractors and Dynamical Regimes in Semantic Space

@@ Abstract
Iterative LLM systems—self-refinement, chain-of-thought, autonomous agents—are increasingly deployed, yet their temporal dynamics remain uncharacterized. Prior work evaluates task performance at convergence but ignores the trajectory: how does semantic content evolve across iterations? Does it stabilize, drift, or oscillate? Without answering these questions, we cannot predict system behavior, guarantee stability, or systematically design iterative architectures.
We formalize agentic loops as discrete dynamical systems in semantic space. Borrowing from dynamical systems theory, we define trajectories, attractors and dynamical regimes for recursive LLM transformations, providing rigorous geometric definitions adapted to this setting. Our framework reveals that agentic loops exhibit classifiable dynamics: contractive (convergence toward stable semantic attractors), oscillatory (cycling among attractors), or exploratory (unbounded divergence).
Experiments on singular loops validate the framework. Iterative paraphrasing produces contractive dynamics with measurable attractor formation and decreasing dispersion. Iterative negation produces exploratory dynamics with no stable structure. Crucially, prompt design directly controls the dynamical regime—the same model exhibits fundamentally different geometric behaviors depending solely on the transformation applied.
This work establishes that iterative LLM dynamics are predictable and controllable, opening new directions for stability analysis, trajectory forecasting, and principled design of composite loops that balance convergence and exploration.
@@ 1 Introduction
Iterative LLM systems are now ubiquitous. Self-Refine Madaan et al. (2023) iteratively critiques and improves outputs. Chain-of-thought Wei et al. (2022) decomposes reasoning into sequential steps. Autonomous agents Yao et al. (2023); Shinn et al. (2023); Park et al. (2023) maintain extended recursive interactions. These systems share a common structure: agentic loops where each output becomes the next input, $a_{t+1}=F(a_{t})$.
Yet despite widespread deployment, a fundamental question remains unanswered: what are the dynamics of these loops? Do they converge to stable outputs? Drift unboundedly? Oscillate between states? Current work evaluates only endpoints—task accuracy after $n$ iterations—while ignoring the trajectory that produced them. This gap is consequential: without understanding dynamics, we cannot predict when loops will stabilize, detect runaway divergence, or design loops with desired convergence properties.
The gap.
Existing research treats iterations as independent steps rather than as a dynamical process. Self-Refine Madaan et al. (2023) measures task improvement but not geometric trajectory structure. Agentic frameworks Yao et al. (2023); Shinn et al. (2023) demonstrate sophisticated behaviors without characterizing convergence or divergence patterns. Embedding geometry research Ethayarajh (2019); Wang and Isola (2020) analyzes static distributional properties but not how representations evolve under repeated transformations.
@@
We formalize agentic loops as discrete dynamical systems in semantic space, drawing on classical theory Strogatz (2018). Our contributions are:
  1. (i)
    Definitions. Operational concepts of trajectory (the sequence of semantic embeddings as the loop evolves), attractor (stable regions detected via incremental clustering), and dynamical regime (contractive, oscillatory, or exploratory). These are tailored to LLM transformations, not generic dynamical systems theory.
  2. (ii)
    Measurement protocol. A concrete pipeline from text to calibrated embedding to geometric indicators. We define and compute local drift (similarity to previous iteration), global drift (similarity to initial state), dispersion, and cluster persistence—all using calibrated similarity Tacheny (2025) to ensure measurements align with human semantic judgments.
  3. (iii)
    Control insight. Empirical demonstration that prompt design determines the dynamical regime. On the same LLM, iterative paraphrasing produces contractive dynamics (similarity increases, dispersion decreases, stable attractors form), while iterative negation produces exploratory dynamics (similarity fluctuates, no stable structure emerges). The transformation operator, not the model, controls the regime.
Why it matters.
These contributions enable capabilities absent from prior work: stability analysis (detect convergence/divergence from early iterations), trajectory prediction (forecast future semantic states), and principled loop design (engineer composite loops with controlled dynamics). For computational creativity Boden (2009); Colton and Wiggins (2012), this provides infrastructure to balance exploration (novelty) and consolidation (value preservation).
Scope.
For improved interpretability, we use a calibrated semantic similarity Tacheny (2025) where measurements align with human judgments. However, our definitions and qualitative results (contraction vs. divergence, attractor detection) are order-based and thus robust to any monotonic transformation of the similarity metric, including raw cosine similarity.